OpenStudy (anonymous):

[02.03] Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

x=-3, x=-11

jigglypuff314 (jigglypuff314):

to solve for x: -3|x + 7| = -12 first divide by -3 on both sides so |x + 7| = 4 which separates out into x + 7 = 4 and -(x+7)=4 now solve for x -7 -7 -x - 7 = 4 x = 4 - 7 x = -4 -7